OpenStudy (anonymous):

Factor the trinomial: 3x^2+7x+2

OpenStudy (anonymous):

youre going to get (x+2)(3x+1)

OpenStudy (anonymous):

Can you explain the steps please

OpenStudy (anonymous):

1. (3x + blank)(x + blank) or (3x - blank)(x - blank)

OpenStudy (anonymous):

look for 2 #'s multeply to get 2 and add up to 7

OpenStudy (anonymous):

since we have 3x to begin with, use some brain sells and you arrive at the answer :)

OpenStudy (anonymous):

(3x + 1)(x + 2)

OpenStudy (anonymous):

remember that factoring is just the undoing of multiplication.

OpenStudy (anonymous):

The only numbers that multiply to get 2 are 1*2 ? and that doesnt add up to 7

OpenStudy (anonymous):

@Vackious ?

OpenStudy (anonymous):

@ineedyouubiebs yo did you read and understand what i posted ?

OpenStudy (mathmate):

Grouping is your best friend! 3x^2+7x+2 we look for two integers m and n such that m*n=3*2=6, and m+n=7 So you would find m=6, n=1 By splitting 7x into 6x+1x, we write the expression as 3x^2+6x + x + 2 =3x(x+2) + (x+2) Factor out the x+2 to get (x+2)(3x+1) exactly as someone posted, but without explanations.